Present and Future Values of Single Cash Flows for Different Periods

Find the following values, using the equations, and then work the problems using a financial calculator to check your answers. Disregard rounding differences. (Hint: If you are using a financial calculator, you can enter the known values and then press the appropriate key to find the unknown variable. Then, without clearing the TVM register, you can "override" the variable that changes by simply entering a new value for it and then pressing the key for the unknown variable to obtain the second answer. This procedure can be used in parts b and d, and in many other situations, to see how changes in input variables affect the output variable.)

  1. An initial $200 compounded for 1 year at 8.9%. Round your answers to the nearest cent.
    $  
  2. An initial $200 compounded for 2 years at 8.9%. Round your answers to the nearest cent.
    $  
  3. The present value of $200 due in 1 year at a discount rate of 8.9%. Round your answers to the nearest cent.
    $  
  4. The present value of $200 due in 2 years at a discount rate of 8.9%. Round your answers to the nearest cent.
    $  
Answer #1

ANSWER:

(a).Future value = Present value (1+interest rate)^n

= 200*(1+8.9%)^1

=200*(1.089)

=$217.8

(b).future value = 200*(1+8.9%)^2

=200*(1.1859)

=$237.184

(c).present value = future value/(1+interest rate)^n

= 200/(1+8.9%)

=200/1.089

=$183.65

(d). Present value = 200/(1+8.9%)^2

=200/1.1859

=$168.64

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************